OpenStudy (anonymous):

What is the value of -(-33)?

OpenStudy (anonymous):



OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

first you evaluate \(\Large 3^3\)

jimthompson5910 (jim_thompson5910):

the cube exponent just tells you to multiply the base by itself 3 times

OpenStudy (anonymous):

So 9

jimthompson5910 (jim_thompson5910):

example: \(\Large 7^3 = 7*7*7 = 343\)

jimthompson5910 (jim_thompson5910):

try again. Look at my example

OpenStudy (anonymous):

27

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

Thats the final answer?

jimthompson5910 (jim_thompson5910):